Given the utility function, U=∂x+ßy, determine the marginal rate of substitution


1
Expert's answer
2021-01-20T12:13:01-0500

Answer:


The given utility function

U = ∂x+βy


We know that marginal rate is given by 


"dy \over dx"= "Ux \over Uy"


U"∂U \over ∂x"  = ∂*1+0= ∂


Uy = "∂U \over ∂y" = 0 + β×1= β


the marginal rate = "∂ \over β"